# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

### v4

@@ 3.2 INT4 Per-thread Quantization
Orthogonal to smoothing, we can mitigate the problem of outliers by refining the quantization granularity so the number of affected elements by outliers becomes smaller. Although per-token offers a detailed level of granularity, it results in significant overhead during dequantization. Specifically, each GPU thread in per-token quantization must handle multiple quantization scales, leading to a high latency of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. SageAttention uses per-block quantization, where each block $Q_{i}$ ($b_{q}$ tokens) and $K_{i}$ ($b_{k}$ tokens) have a single quantization scale. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the high dequantization overhead. However, quantizing $Q$ and $K$ to INT4 demands a finer quantization granularity. To address this, we propose per-thread quantization, a more precise and granular approach than the per-block quantizer, also without the additional overhead of vectors dot product between $\delta_{Q}$ and $\delta_{K}$.
Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM). We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ since $K_{j}$ is shared among warps. Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$. According to the layout requirement of this instruction, we find that $Q_{w}[8k+i]$ could share one quantization scale, and $K_{j}[8k+2i]$ and $K_{j}[8k+2i+1]$ could share one quantization scale. Such a quantization method is more fine-grained and with no additional overhead. This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value. We show an example of per-thread quantization in Fig. 4. The detailed formulation is shown in Equation A.5 and Fig. 17 (Please see Appendix A.5 for more detail).
Empirical results. As shown in Table 6 and Table 15, we compare the average and the worst accuracy of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate the accuracy of per-thread quantization is very close to per-token and outperforms a lot more than others.
@@
  Preprocessing: $K=K-\mathrm{mean}(K)$, ${}~{}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. //per-channel.
@@
       $m_{ij}=\mathrm{max}(m_{i,j-1},\mathrm{rowmax}(S_{ij}))$, ${}\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$, ${}l_{ij}=e^{m_{i,j-1}-m_{ij}}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
@@
[…]
@@ 3.3 FP8 quantization for $\tilde{P}V$
We now turn to the MM $\tilde{P}V$, where $\tilde{P}_{ij}=\exp(S_{ij}-m_{ij})$ is the unnormalized quantity according to Eq. (1). The distribution of $\tilde{P}$ is unique and differs from other activations. First, we note that $S_{ij}-m_{ij}\leq 0$, so $P_{ij}\in[0,1]$ ($\leq$ and $\in$ apply element-wisely). We find that $\tilde{P}$ often consists of many small elements, but their sum is non-negligible (e.g., 5000 elements around $10^{-4}$). In this case, we must represent small elements accurately. INT quantization is unsuitable for this proposal since it puts the quantization points evenly within the numerical range. SageAttention (Zhang et al., 2025) choose to retain $\tilde{P}$ and $V$ in FP16, and accelerate the MM by decreasing the accumulator precision. However, this strategy is only effective on very few GPUs.
We propose to quantize $\widetilde{P},V$ to FP8 with 4 exponent bits and 3 mantissa bits (E4M3). The numerical range of E4M3 is $[-448,+448]$. We quantize $P$ with a static scale: $\delta_{P}=\frac{1}{448}$ since the original $P$ elements is already in $[0,1]$. We quantize $V$ per-channel to address the channel-wise outliers shown in Fig. 3. Empirical results in Table 7 and Table 16 show the average and the worst accuracy of different data types used for $\widetilde{P},V$ across all layers of CogvideoX. The accumulator is always 32-bit. We can see that the accuracy of E4M3 is very close to that of FP16 and superior to E5M2 and INT8. Most modern GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16.
@@ 3.4 FP32 MMA Buffer for FP22 Accumulator
While FP8 quantization for $\tilde{P}V$ above is theoretically accurate in simulation, we observe that the actual CUDA implementation suffers a consistent accuracy degradation. After narrowing down the problem, we find that the accumulator for the mma(f32f8f8f32) instruction on the Ada and Hopper architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are FP8 matrices and $C,D$ are FP32 matrices, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. When $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ matches the result of the high-precision mma(f16f16f32) instruction. However, when $D$ is initialized with more than 13 mantissa bits, the error of $C$ corresponds to the difference between the results of mma(f32f16f16f32) and mma(f32f8f8f32). Consequently, the matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
@@
Remark. The two-level accumulation strategy is also implemented in Cutlass (NVIDIA, 2023) and DeepSeekv3 (DeepSeek-AI et al., 2024) for computing activation-weight products in linear layers. To the best of our knowledge, we are the first to discover and investigate the effect of the FP22 accumulator and implement the two-level accumulation for attention.
Optional smooth V technique. We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases: $\overrightarrow{V_{m}}=\mathrm{mean}(V,\text{axis}=0),V=V-\overrightarrow{V_{m}}$. Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
@@ 4 Experiment
Main result. SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs and is much more accurate than FlashAttention3(fp8). Our approach maintains end-to-end metrics across language, image, and video generation models.
@@ 4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2025), HunyuanVideo (Kong et al., 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
@@
Implemetation. We implement two attention kernels as shown in Table 3 using CUDA.
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2024), we apply smooth quant for $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for $Q,K$ before INT4 quantization. (3) SageAttention (Zhang et al., 2025), which uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$. (4) FlashAttn3(fp8), the FP8 version of FlashAttention3, which only runs on Hopper GPUs.
@@
Model GPU Original SageAttn 2-8b SageAttn 2-4b
@@ 4.2 Speed and Accuracy of Kernels
Kernel Speed. We compare the Speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017). Specifically, Fig. 5 shows the speed across varying sequence lengths on RTX4090, indicating that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively. Fig. 91011121314, and 15 in Appendix A.2 show more kernel speed comparison on RTX4090, L20, H20, H100 GPUs.
@@ 4.3 End-to-end Performance
Visible image and video examples. Fig.6 and Fig. 8 in Appendix A.1 show some visible comparision examples from HunyuanVideo, Mochi and CogvideoX (1.5-5B). We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences but is much better than the baselines.
@@ 5 Conclusion
We introduce SageAttention2, an efficient and accurate quantized attention. First, we propose to quantize matrixes $(Q,K)$ in a thread-level granularity and $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of $QK^{\top}$. Third, we propose a two-level accumulation strategy to enhance the accuracy of FP8 $\widetilde{P}V$. SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, but offers significantly higher accuracy. Extensive testing confirms that our approach maintains end-to-end metrics across language, image, and video generation models.
@@ A.1 Visible Comparison Exmaples
Figure 8: A comparison example using SageAttn2-8b and FlashAttention3 on CogvideoX-1.5, Mochi and HunyuanVideo.
@@ A.2 Additional Kernel Speed Comparison
Fig. 9,  10,  11,  12,  13,  14, and  15 compare the speed of SageAttention2 against baselines using configurations with headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017), on RTX4090, L20, H100, and H20 GPUs.
@@ A.3 Smoothing V
As shown in Fig. 16, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ in some models consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9 in CogvideoX. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean $\overrightarrow{V_{m}}$ along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in higher representational precision (see Fig. 16 for a visual demonstration). Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.

### v7

@@ 3.2 INT4 Per-thread Quantization
Orthogonal to smoothing, we can mitigate the problem of outliers by refining the quantization granularity so that the number of affected elements by outliers becomes smaller. Although per-token quantization offers a detailed level of granularity, it results in significant overhead during dequantization. Specifically, each GPU thread in per-token quantization must handle multiple quantization scales, leading to a high latency of the dot product of the quantization scale vectors $\delta_{Q}$ and $\delta_{K}$. SageAttention uses per-block quantization, where each block $Q_{i}$ ($b_{q}$ tokens) and $K_{i}$ ($b_{k}$ tokens) have a single quantization scale. Such a quantization strategy could achieve an accuracy performance close to per-token quantization and avoid the high dequantization overhead. However, quantizing $Q$ and $K$ to INT4 demands a finer quantization granularity. To address this, we propose per-thread quantization, a more precise and granular approach than the per-block quantizer, also without the additional overhead of the vector dot product between $\delta_{Q}$ and $\delta_{K}$.
Specifically, each block of $Q$, i.e., $Q_{i}$, in SageAttention will be split into $c_{w}$ segments and processed by $c_{w}$ GPU warps in a GPU streaming processor (SM). We call each segment of $Q_{i}$ as $Q_{w}$, and $k_{w}=K_{j}$ since $K_{j}$ is shared among warps. Then, each warp containing 32 threads uses the mma.m16n8k64 PTX instruction (NVIDIA, ) for the $Q_{w}K_{j}^{\top}$. According to the layout requirement of this instruction, we find that $Q_{w}[8k+i]$ could share one quantization scale, and $K_{j}[8k+2i]$ together with $K_{j}[8k+2i+1]$ could share one quantization scale. Such a quantization method is more fine-grained with no additional overhead. This is because it assigns different GPU threads to distinct quantization groups based on the MMA instruction layout, with each thread performing dequantization only using a single quantization scale value. We show an example of per-thread quantization in Fig. 4. The detailed formulation is shown in Equation A.6 and Fig. 18 (please see Appendix A.6 for more detail).
Empirical results. As shown in Table 6 and Table 15, we compare the average and worst accuracy of INT4 quantization at per-token, per-thread, per-block, and per-tensor granularity using real $Q,K,V$ across all layers of CogvideoX. Results indicate that the accuracy of per-thread quantization is very close to per-token and significantly outperforms other granularities. Moreover, Table 19 shows that per-thread quantization incurs almost no speed degradation, while per-token quantization introduces noticeable overhead due to the reduced hardware efficiency.
@@
  Preprocessing: $K=K-\mathrm{mean}(K)$, ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{(\delta_{V},\hat{V})}}={\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\psi_{V}}}(V)$. //per-channel.
@@
   $m_{ij}=\mathrm{max}(m_{i,j-1},\mathrm{rowmax}(S_{ij}))$, $\widetilde{P}_{ij}=\mathrm{exp}(S_{ij}-m_{ij})$, $l_{ij}=e^{m_{i,j-1}-m_{ij}}l_{i,j-1}+\mathrm{rowsum}(\widetilde{P}_{ij})$ ;
@@
  Load $\delta_{V}$ into an SM ;    $O_{i}=\mathrm{diag}(l_{i,T_{n}})^{-1}O_{i,T_{n}}\text{(FP32)}$ $448*\delta_{V}$ ;    Write $O_{i}$ ;
@@ 3.3 FP8 quantization for $\tilde{P}V$
We now turn to the MM $\tilde{P}V$, where $\tilde{P}_{ij}=\exp(S_{ij}-m_{ij})$ is the unnormalized quantity according to Eq. (1). The distribution of $\tilde{P}$ is unique and differs from other activations. First, we note that $S_{ij}-m_{ij}\leq 0$, so $P_{ij}\in[0,1]$ ($\leq$ and $\in$ apply element-wise). We find that $\tilde{P}$ often consists of many small elements, but their sum is non-negligible (e.g., 5000 elements around $10^{-4}$). In this case, we must represent small elements accurately. INT quantization is unsuitable for this setting, since it distributes the quantization points evenly within the numerical range. SageAttention (Zhang et al., 2025c) choose to retain $\tilde{P}$ and $V$ in FP16, and accelerate the MM by decreasing the accumulator precision. However, this strategy is only effective on very few GPUs.
We propose to quantize $\widetilde{P},V$ to FP8 with 4 exponent bits and 3 mantissa bits (E4M3). The numerical range of E4M3 is [448,+448][-448,+448]. We quantize $P$ with a static scale: $\delta_{P}=\frac{1}{448}$ since the original $P$ elements are already in $[0,1]$. We quantize $V$ per-channel to address the channel-wise outliers shown in Fig. 2. Empirical results in Table 7 and Table 16 show the average and worst accuracy of different data types used for $\widetilde{P},V$ across all layers of CogvideoX. The accumulator is always 32-bit. We can see that the accuracy of E4M3 is very close to that of FP16 and superior to E5M2 and INT8. Most modern GPUs have tensor cores that support FP8 Matmul operations, which are twice as fast as those using FP16.
@@ 3.4 FP32 MMA Buffer for FP22 Accumulator
While FP8 quantization for $\tilde{P}V$ above is theoretically accurate in simulation, we observe that the actual CUDA implementation suffers a consistent accuracy degradation. After narrowing down the problem, we find that the accumulator for the mma(f32f8f8f32) instruction on the Ada and Hopper architecture is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. Specifically, for mma(f32f8f8f32) instruction $C=AB+D$, where $A,B$ are FP8 matrices and $C,D$ are FP32 matrices, we initialize the $A,B$ to zero and vary $D$ to test the data type of the accumulator. When $D$ is initialized with 1 sign bit, 8 exponent bits, and 13 mantissa bits, the value of $C$ exactly matches $D$. However, when $D$ is initialized with more than 13 mantissa bits, the value of $C$ is equal to $D$ with its least significant 10 mantissa bits zeroed out (i.e., truncated). Consequently, matrix multiplication of $\widetilde{P}V$, quantized to FP8, incurs a certain degree of accuracy loss compared to using an FP32 accumulator.
@@
Remark. The two-level accumulation strategy is also implemented in CUTLASS (NVIDIA, 2023) and DeepGemm (DeepSeek-AI et al., 2024) for computing weight-activation products in linear layers. To the best of our knowledge, we are the first to discover and investigate the effect of the FP22 accumulator and implement the two-level accumulation for attention.
Optional smooth V technique. We also figure out another way to mitigate the accuracy loss due to the FP22 accumulator when $V$ possesses channel-wise biases: $\overrightarrow{V_{m}}=\mathrm{mean}(V,\text{axis}=0),V=V-\overrightarrow{V_{m}}$. Furthermore, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$.
@@
Figure 7: A comparison example using SageAttn2-8b and FlashAttention3 on CogvideoX-1.5.
@@ 4 Experiment
Main result. SageAttention2 is faster than FlashAttention2 and xformers by about 3x and 4.5x. Moreover, SageAttention2 matches the speed of FlashAttn3(fp8) on the Hopper GPUs and is much more accurate than FlashAttn3(fp8). SageAttention2 maintains end-to-end metrics across language, image, and video generation models.
@@ 4.1 Setup
Models. We validate the effectiveness of SageAttention2 across a diverse set of representative models from language, image, and video generation. Specifically, we conduct experiments on ten models: Llama2 (7B) (Touvron et al., 2023), Llama3.1 (8B) (Dubey et al., 2024), and GLM4 (9B) (GLM et al., 2024) for text2text, CogvideoX (2B), CogvideoX (1.5-5B) (Yang et al., 2025b), HunyuanVideo (Kong et al., 2024), and Mochi (Team, 2024) for text2video, Flux (schnell) (Black Forest Labs, 2023) and Stable-Diffusion3.5 (turbo) (Stability AI, 2023) for text2image, and TIMM (Wightman, 2019) for image classification.
@@
Implemetation. We implement two attention kernels as shown in Table 3 using CUDA. The 8-bit variant is adapted for NVIDIA Hopper GPUs, which lack native INT4 tensor core support, and incorporates all techniques described in Sec. 3 except for smoothing Q.
Baselines. (1) SmoothAttn. Following Qserve (Lin et al., 2025), we apply smooth quant for $Q,K$ with smoothing factor $\alpha=0.5$. (2) HadmdAttn. Following Quarot (Ashkboos et al., 2024), we apply random Hadamard transformation for $Q,K$ before INT4 quantization. (3) SageAttention (Zhang et al., 2025c), which uses smoothing $K$, INT8 per-block quantization for $Q,K$, and FP16 for $\widetilde{P},V$. (4) FlashAttn3(fp8), the FP8 version of FlashAttention3, which only runs on Hopper GPUs.
@@
Model GPU Original
SageAttn
2-8b
SageAttn
2-4b
@@ 4.2 Speed and Accuracy of Kernels
Kernel Speed. We compare the speed of SageAttention2 against baselines using headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017). Detailed setup can be found in Appendix A.8. Specifically, Fig. 5 shows the speed across varying sequence lengths on RTX4090, indicating that SageAttn2-4b and SageAttn2-8b are approximately 3x and 2.7x faster than FlashAttention2, and about 4.5x and 4x faster than xformers, respectively. Fig. 101112131415, and 16 in Appendix A.2 show more speed results on RTX4090, L20, H20, H100 GPUs.
@@ 4.3 End-to-end Performance
Visible image and video examples. Fig. 678, and 9 show some visible comparison examples from HunyuanVideo, Mochi and CogvideoX. We can observe that SageAttn2-8b does not introduce any visible differences compared to full-precision attention, whereas SageAttn2-4b has minor differences but is much better than the baselines.
@@ 5 Conclusion
We introduce SageAttention2, an efficient and accurate quantized attention. First, we propose to quantize matrices $(Q,K)$ in a thread-level granularity and $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of $QK^{\top}$. Third, we propose a two-level accumulation strategy to enhance the accuracy of FP8 $\widetilde{P}V$. SageAttention2 is faster than FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, but offers significantly higher accuracy. Extensive experiments confirm that our approach maintains end-to-end metrics across language, image, and video generation models.
Acknowledgment
This work was supported by the NSFC Projects (Nos. 92270001, 62376131). J.Z is also supported by the XPlorer Prize.
Impact Statement
This paper presents work that aims to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Visible Comparison Exmaples
Figure 9: A comparison example using SageAttn2-8b and FlashAttention3 on Mochi and HunyuanVideo.
@@ A.2 Additional Kernel Speed Comparison
Fig. 10,  11,  12,  13,  14,  15, and  16 compare the speed of SageAttention2 against baselines using headdim=64 and headdim=128, both with and without Causal Mask (Vaswani, 2017), on RTX4090, L20, H100, and H20 GPUs.
@@ A.3 Smoothing V
As shown in Fig. 17, this strategy could enhance the accuracy of FP22 for values in $\widetilde{P}V$ for the following reasons: Each row of $\widetilde{P}$ spans a value range from 0 to 1, and each column of $V$ in some models consistently features channel-wise biases that are exclusively positive or negative, for instance, ranging between 8 and 9 in CogvideoX. Consequently, the values of $\widetilde{P}V$ could be quite large. However, the floating-point number representation range is not uniform—it is denser near zero. Therefore, by subtracting the mean $\overrightarrow{V_{m}}$ along the channel dimension from $V$, the values of $\widetilde{P}V$ will be closer to zero, resulting in a higher representational precision (see Fig. 17 for a visual demonstration). Additionally, to maintain the correctness of the attention computation, it is only necessary to add $\overrightarrow{V_{m}}$ to the final calculation of $O$: $O=O+\overrightarrow{V_{m}}$. This is because the sum of each row of the $\widetilde{P}$ matrix equals 1, so $\widetilde{P}\overrightarrow{V_{m}}=\overrightarrow{V_{m}}$. In other words, this method decomposes $V$ into two parts: $\overrightarrow{V_{m}}$ and $V$. For $V$, it centers the values of each column around zero, which leads to the dot product result between a row from the quantized $\widetilde{P}$ matrix and a column from the quantized $V$ matrix being closer to zero. This makes the representation of FP22 more accurate. Meanwhile, $\overrightarrow{V_{m}}$ is retained in FP16 and is added to $O$ at the end without causing a loss of precision for the $\overrightarrow{V_{m}}$ part.